# Learning Large-Scale Competitive Team Behaviors with Mean-Field Interactions and Online Opponent Modeling

### v2

@@ 1 Introduction
Existing state-of-the-art MARL algorithms built upon MADDPG, MAAC and MA-PPO [Lowe et al.,, 2017; Yu et al.,, 2022], face severe scalability challenges as the number of agents grows, primarily due to the well-known curse of dimensionality. A promising remedy is offered by mean-field theory, which approximates large-scale agent–environment interactions in the infinite-population limit [Huang et al.,, 2006]. Two major areas of mean-field research are mean-field games (MFGs) [Huang et al.,, 2006; Lasry and Lions,, 2007; Sen and Caines,, 2019; Laurière et al.,, 2022], which focus on non-cooperative agents, and mean-field control (MFC) problems [Bensoussan et al.,, 2013; Gu et al.,, 2021], which study fully cooperative scenarios. In contrast, mixed collaborative–competitive scenarios that arise in many real-world domains, such as team sports [Gaviria Alzate et al.,, 2025] and social dilemmas [Leibo et al.,, 2017], remain relatively underexplored. To address this gap, we propose Mean-Field Multi-Agent Proximal Policy Optimization (MF-MAPPO), the first PPO-based learning algorithm tailored for mixed cooperative–competitive mean-field settings. Guided by existing theoretical results [Guan et al., 2024b, ], MF-MAPPO scales to hundreds or thousands of agents while preserving convergence guarantees and remaining agnostic to individual identities or private observations.
Mean-Field Teams. The single-team problem was explored in Arabneydi and Mahajan, [2014], where agents share a common team reward (MFC). In contrast, Mahajan and Nayyar, [2015] established optimality for finite-population games only in the LQG setting, while Sanjari et al., [2022] analyzed a two-team setting with continuous states and actions—unlike our finite state-space formulation that directly admits the familiar MDP-type structure. Similarly, multi-population MFGs (MP-MFGs) have been studied in the past but often restrict agent dynamics and policies to be independent of both other agents and other population distributions, see Perolat et al., [2021] and references therein. More recently, Guan et al., 2024b introduced zero-sum mean-field team games (ZS-MFTGs), modeling large-population teams that compete while cooperating internally. A common-information decomposition [Nayyar et al.,, 2013] enables team-size-independent learning with identical optimal team policies using MF feedback, unlike aforementioned MP-MFGs (open-loop MF policies). However, computing such policies numerically, especially in large state-action spaces using dynamic programming, is costly. By contrast, MF-MAPPO leverages shared actor–critic networks per team and uses only commonly accessible information (exact/estimated), ensuring tractability and scalability.
@@
Mean-Field Theory and Learning. Recent advances in mean-field learning span Q-function–based methods, such as MF-Q and MF-AC [Yang et al.,, 2020] and DDPG-MFTG [Shao et al., 2024b, ], to value-function–based methods like Dec-POMFPPO [Cui et al.,, 2024] (MFC only). While DDPG-MFTG incorporates team games, it is restricted to simple grid worlds, unlike our focus on tightly coupled collaborative–competitive domains (ZS-MFTGs). We adopt it as a baseline and show that MF-MAPPO consistently outperforms it in stability and performance. Other related works include PMD-TD for MFGs [Yardim and He,, 2024] and GAN-based ECA-Net for continuous-space attack–defense games [Wang et al.,, 2022], both differing in scope and structure. Moreover, Yang et al., [2020] define mean fields over neighboring actions rather than the full state space. MF-MAPPO extends PPO [Schulman et al.,, 2017] to competitive MFTGs, using team distributions as critic inputs for scalability, a shared actor–critic per team with a single buffer for efficiency, and simultaneous team training to avoid the inefficiencies of iterative best-response methods [Lanctot et al.,, 2017; Smith et al.,, 2021]. Unlike prior MF methods that rely on infinite-population oracles [Shao et al., 2024a, ; Perolat et al.,, 2021; Carmona et al.,, 2021], MF-MAPPO is trained directly in finite-population simulators, making it suitable for realistic deployment. Finally, to standardize evaluation in large-scale MFTGs, we create novel MFTG benchmark environments (Constrained Rock–Paper–Scissors, Battlefield) going beyond existing ones that either focus only on MFGs [Guo et al.,, 2023] or omit MF coupling altogether [Terry et al.,, 2021; Zheng et al.,, 2018].
@@
Mean-Field Estimation and Opponent Modeling. To enable reactive behavior to opponents’ actions, most existing MF approaches assume centralized or exact knowledge of the opponent’s MF, which is rarely practical. While Cui et al., [2024] considers partial observability, it is limited to MFC. Estimation methods such as kernel density estimation [Inoue et al.,, 2021] or normalizing flows [Perrin et al.,, 2021] struggle with discrete spaces and limited agent-visibility. Communication-based methods [Benjamin and Abate, 2025b, ; Benjamin and Abate, 2025a, ] address these constraints but assume uniform estimates for unobserved states and rely on perfect multi-round communication while being restricted to fully cooperative or competitive regimes; noisy variants may even produce invalid distributions. We instead propose Dynamic-Projected Consensus (D-PC), a constrained consensus algorithm that ensures validity, exponential convergence, and bounded deviations when paired with a gradient-regularized MF-MAPPO policy. Gradient regularization naturally stabilizes MF-MAPPO in partially observable MFTGs. Experiments show that D-PC matches baseline performance and even outperforms them under limited communication, especially critical in adversarial settings requiring rapid adaptation [Richards et al.,, 2012], enhancing robustness and fault tolerance. To our knowledge, this is the first use of MF estimation for opponent modeling in competitive team settings.
@@ 2.1 Zero-Sum Mean-Field Team Game
The zero-sum mean-field team game models a discrete-time stochastic game between two large teams of agents [Guan et al., 2024b, ]. The Blue and Red teams consist of $N_{1}$ and $N_{2}$ identical agents for each team, with the total number of agents being $N\hskip-1.4457pt=\hskip-1.4457ptN_{1}\hskip-1.4457pt+\hskip-1.4457ptN_{2}$. Let $X^{N_{1}}_{i,t}\in\mathcal{X}$ and $U^{N_{1}}_{i,t}\in\mathcal{U}$ represent the state and action of Blue agent $i\in[N_{1}]$ at time $t$. Here, $\mathcal{X}$ and $\mathcal{U}$ are the finite state and action spaces of the Blue team. Similarly, $Y^{N_{2}}_{j,t}\in\mathcal{Y}$ and $V^{N_{2}}_{j,t}\in\mathcal{V}$ denote the state and action of Red agent $j\in[N_{2}]$ at time $t$. The joint state-action variables for the Blue and Red teams are denoted as $(\mathbf{X}^{N_{1}}_{t},\mathbf{U}^{N_{1}}_{t})$ and $(\mathbf{Y}^{N_{2}}_{t},\mathbf{V}^{N_{2}}_{t})$, respectively. We denote the space of probability measures over a set $E$ as $\mathcal{P}({E})$. Below, $\mathrm{d}_{\mathrm{TV}}\big(\mu,\mu^{\prime}\big)$ represents the total variation between $\mu,\mu^{\prime}\in\mathcal{P}(E)$.
@@
We consider weakly-coupled dynamics where the dynamics of each individual agent is coupled with other agents through the EDs [Huang et al.,, 2006; Sanjari et al.,, 2022]. For Blue agent $i$, its stochastic transition is governed by the transition kernel $f_{t}:\mathcal{X}\times\mathcal{U}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\rightarrow\mathcal{P}(\mathcal{X})$:
@@
Lipschitz continuity is commonly assumed [Huang et al.,, 2006; Gu et al.,, 2021], and at minimum uniform continuity is required; see Cui et al., [2024] for counterexamples.
@@
We consider a mean-field sharing information structure [Arabneydi and Mahajan,, 2015], where each agent’s decision depends on its own state and the two team EDs. We start with assuming full observation of mean-fields and later relax this assumption. Specifically, the Blue and Red agents seek to construct mixed Markov policies $\phi_{i,t}:\mathcal{U}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\to[0,1],~~\text{and}~~\psi_{j,t}:\mathcal{V}\times\mathcal{Y}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\to[0,1]$, where the Blue policy $\phi_{i,t}(u|x_{i,t}^{N_{1}},\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ dictates the probability that Blue agent $i$ selects action $u$ given its state $x_{i,t}^{N_{1}}$ and the observed/estimated team EDs $\mu^{N_{1}}_{t}$ and $\nu^{N_{2}}_{t}$. Note that each agent’s individual state is its private information.
@@ 2.2 Infinite-Population Solution
To reduce the complexity of team policy optimization domains in (4), the authors of Guan et al., 2024b examined team behaviors under identical team policies at the infinite-population limit. It was shown that the team joint states can be represented using the team population distribution, which coincides with the state distribution of a typical agent referred to as the mean-fields ($\mu_{t}$ and $\nu_{t}$ for the Blue and Red teams, respectively). They also proved that MFs induced by identical team policies in an infinite-population game closely approximate the EDs induced by non-identical team policies in the corresponding finite-population game, which justifies the simplification of the optimization domain in (4) to identical team policies (also see Theorem 1). Furthermore, there is a one-to-one correspondence between infinite-population coordination policies $(\alpha,\beta)$ and local identical team policies $(\phi,\psi)\in\Phi\times\Psi$. The performance of $(\phi,\psi)$ in the equivalent zero-sum coordinator game is measured by
@@
where $\mu_{t}$ and $\nu_{t}$ follow a deterministic dynamics [Guan et al., 2024b, ] similar to the state distribution propagation of a controlled Markov chain. The worst-case performance of the Blue team in this infinite-population game is then given by the lower game value $\underline{J}_{\infty}^{*}(\mu_{0},\nu_{0})=\max_{\phi\in\Phi}\;\min_{\psi\in\Psi}~J_{\infty}^{\phi,\psi}(\mu_{0},\nu_{0})$, where the optimization domain is restricted to identical team policies. Guan et al., 2024b establishes guarantees that identical team policies resulting from the solution of this equivalent zero-sum coordinator game are still $\epsilon$-optimal for the original max-min optimization problem in (4) where $\epsilon=\mathcal{O}(1/\sqrt{\underline{N}})$ and $\underline{N}=\min\{N_{1},N_{2}\}$.
The infinite-population limit of large-population games offers several theoretical advantages, such as representing the population by a typical agent and deterministic dynamics that reduce (3) to the non-stochastic optimization of (5). Previous works [Shao et al., 2024a, ; Perolat et al.,, 2021; Carmona et al.,, 2021] depend on infinite-population oracles to obtain mean-field trajectories $(\mu_{t},\nu_{t})$ in order to compute (5). This is rather unrealistic, since in practice, only finite-population simulations and local states $(\mathbf{x}^{N_{1}}_{t},\mathbf{y}^{N_{2}}_{t})$ with actions $(\mathbf{u}^{N_{1}}_{t},\mathbf{v}^{N_{2}}_{t})$ are available/observable. Moreover, a single coordinator policy $\alpha(\beta)$ defines a distribution over actions for each state conditioned on the mean-field, causing its dimensionality to scale with the joint state–action space (e.g., DDPG-MFTG), leading to high computational cost and degraded empirical performance (see Section 5). In summary, the infinite-population model is both impractical (due to oracle dependency) and computationally intractable (due to policy size). Thus, we turn to finite-population simulators and derive guarantees of optimality, scalability, and convergence of the policy gradient to the infinite-population ZS-MFTG. The next result quantifies the level of suboptimality for the Blue team when it deploys the optimal identical policy learned directly from the solution of finite-population ZS-MFTG.
@@
Theorem 1 provides a principled justification for learning identical finite-population team policies in competitive–collaborative team games even when being exploited by non-identical opponent team strategies. Its motivation and proof build on the performance guarantees of the ZS-MFTG in the infinite-population limit, i.e., the coordinator game. Moreover, the error vanishes as $N_{1},N_{2}\to\infty$, thereby recovering the well-studied infinite-population MF formulation [Huang et al.,, 2006]. We detail this finite-population training paradigm in the next section.
@@ 3 Mean-Field Multi-Agent Proximal Policy Optimization
Motivated by Theorem 1, we present an algorithm to learn the finite-population optimal identical team policy. We build our algorithm based on the proximal policy optimization (PPO) framework due to its simplicity and effectiveness. While PPO has shown promising performance in cooperative tasks including MFC problems [Yu et al.,, 2022; Cui et al.,, 2024], its application in mixed competitive-collaborative scenarios is less studied. In the sequel, we introduce our key contribution: MF-MAPPO. We initialize two pairs of actor-critic networks, one for each team, deployed to learn the identical policy used by each team, see Figure 1(b). Specifically, we introduce a minimally-informed critic network by exploiting the MF information structure. The key point here is that we only require commonly accessible information for the critic network in order to learn the value function (Proposition 1). Further, the private information available to each agent only individually enters the actor during training. This results in neural networks that scale well with the number of agents. We present the team actor-critic networks from the Blue team’s perspective, and due to symmetry results extend naturally to the Red team.
@@
Shared-Team Actor. As discussed in earlier sections, the coordinator game is a useful theoretical construct but has limited practical value for real-world deployment since it relies on an infinite-population oracle for training and produces policies whose size scales poorly with the state-action space. We therefore directly train finite-population identical local policies, which preserve the mean-field structure while reducing complexity and improving tractability, with guarantees in Theorem 1. Not only is the approach computationally tractable in terms of the size of the policy, but is also more realistic in terms of sampling training data. We use a single actor network per team to learn identical team policies. The actor optimizes a PPO-based objective with a decaying entropy bonus [Schulman et al.,, 2017; Huang et al.,, 2022], which promotes exploration and stabilizes learning in mean-field settings [Cui and Koeppl,, 2022; Guan et al.,, 2022]. Permutation invariance and identical team policies further allow a single replay buffer per team, reducing memory costs and simplifying experience collection. The PPO-based objective function of the Blue actor is given by:
@@
where, $g(\theta)=\frac{\phi_{\theta}(u|x,\mu,\nu)}{\phi_{\theta^{\text{old}}}(u|x,\mu,\nu)}$, $A_{k}$ is the generalized advantage function estimate [Schulman et al.,, 2018] and the tunable parameter $\omega$ weighs the contribution of the entropy term $S\big(\phi_{\theta}(x,\mu,\nu)\big)$ and decays as training progresses.

### v3

@@ 1 Introduction
Existing state-of-the-art MARL algorithms built upon MADDPG, MAAC and MA-PPO [Lowe et al., 2017; Yu et al., 2022], face severe scalability challenges as the number of agents grows, primarily due to the well-known curse of dimensionality. A promising remedy is offered by mean-field theory, which approximates large-scale agent–environment interactions in the infinite-population limit [Huang et al., 2006]. Two major areas of mean-field research are mean-field games (MFGs) [Huang et al., 2006; Lasry and Lions, 2007; Sen and Caines, 2019; laurière2022scalabledeepreinforcementlearning], which focus on non-cooperative agents, and mean-field control (MFC) problems [Bensoussan et al., 2013; Gu et al., 2021], which study fully cooperative scenarios. In contrast, mixed collaborative–competitive scenarios that arise in many real-world domains, such as team sports [Gaviria Alzate et al., 2025] and social dilemmas [Leibo et al., 2017], remain relatively underexplored. To address this gap, we propose Mean-Field Multi-Agent Proximal Policy Optimization (MF-MAPPO), the first PPO-based learning algorithm tailored for mixed cooperative–competitive mean-field settings. Guided by existing theoretical results [Guan et al., 2024b], MF-MAPPO scales to hundreds or thousands of agents while preserving convergence guarantees and remaining agnostic to individual identities or private observations.
Mean-Field Teams. The single-team problem was explored in Arabneydi and Mahajan [2014], where agents share a common team reward (MFC). In contrast, Mahajan and Nayyar [2015] established optimality for finite-population games only in the LQG setting, while Sanjari et al. [2022] analyzed a two-team setting with continuous states and actions—unlike our finite state-space formulation that directly admits the familiar MDP-type structure. Similarly, multi-population MFGs (MP-MFGs) have been studied in the past but often restrict agent dynamics and policies to be independent of both other agents and other population distributions, see Perolat et al. [2021] and references therein. More recently, Guan et al. [2024b] introduced zero-sum mean-field team games (ZS-MFTGs), modeling large-population teams that compete while cooperating internally. A common-information decomposition [Nayyar et al., 2013] enables team-size-independent learning with identical optimal team policies using MF feedback, unlike aforementioned MP-MFGs (open-loop MF policies). However, computing such policies numerically, especially in large state-action spaces using dynamic programming, is costly. By contrast, MF-MAPPO leverages shared actor–critic networks per team and uses only commonly accessible information (exact/estimated), ensuring tractability and scalability.
@@
Mean-Field Theory and Learning. Recent advances in mean-field learning span Q-function–based methods, such as MF-Q and MF-AC [Yang et al., 2020] and DDPG-MFTG [Shao et al., 2024b], to value-function–based methods like Dec-POMFPPO [Cui et al., 2024] (MFC only). While DDPG-MFTG incorporates team games, it is restricted to simple grid worlds, unlike our focus on tightly coupled collaborative–competitive domains (ZS-MFTGs). We adopt it as a baseline and show that MF-MAPPO consistently outperforms it in stability and performance. Other related works include PMD-TD for MFGs [Yardim and He, 2024] and GAN-based ECA-Net for continuous-space attack–defense games [Wang et al., 2022], both differing in scope and structure. Moreover, Yang et al. [2020]; Subramanian et al. [2020, 2022a, 2022b] define mean fields are defined over (neighboring) agents’ actions rather than their state distribution yielding a rich line of research that is orthogonal to our formulation. MF-MAPPO extends PPO [Schulman et al., 2017] to competitive MFTGs, using team distributions as critic inputs for scalability, a shared actor–critic per team with a single buffer for efficiency, and simultaneous team training to avoid the inefficiencies of iterative best-response methods [Lanctot et al., 2017; Smith et al., 2021]. Unlike prior MF methods that rely on infinite-population oracles [Shao et al., 2024a; Perolat et al., 2021; Carmona et al., 2021], MF-MAPPO is trained directly in finite-population simulators, making it suitable for realistic deployment. Finally, to standardize evaluation in large-scale MFTGs, we create novel MFTG benchmark environments (Constrained Rock–Paper–Scissors, Battlefield) going beyond existing ones that either focus only on MFGs [Guo et al., 2023] or omit MF coupling altogether [Terry et al., 2021; Zheng et al., 2018].
@@
Mean-Field Estimation and Opponent Modeling. To enable reactive behavior to opponents’ actions, most existing MF approaches assume centralized or exact knowledge of the opponent’s MF, which is rarely practical. While Cui et al. [2024] considers partial observability, it is limited to MFC. Estimation methods such as kernel density estimation [Inoue et al., 2021] or normalizing flows [Perrin et al., 2021] struggle with discrete spaces and limited agent-visibility. Communication-based methods [Benjamin and Abate, 2025b, a] address these constraints but assume uniform estimates for unobserved states and rely on perfect multi-round communication while being restricted to fully cooperative or competitive regimes; noisy variants may even produce invalid distributions. We instead propose Dynamic-Projected Consensus (D-PC), a constrained consensus algorithm that ensures validity, exponential convergence, and bounded deviations when paired with a gradient-regularized MF-MAPPO policy. Gradient regularization naturally stabilizes MF-MAPPO in partially observable MFTGs. Experiments show that D-PC matches baseline performance and even outperforms them under limited communication, especially critical in adversarial settings requiring rapid adaptation [Richards et al., 2012], enhancing robustness and fault tolerance. To our knowledge, this is the first use of MF estimation for opponent modeling in competitive team settings.
@@ 2.1 Zero-Sum Mean-Field Team Game
The zero-sum mean-field team game models a discrete-time stochastic game between two large teams of agents [Guan et al., 2024b]. The Blue and Red teams consist of $N_{1}$ and $N_{2}$ identical agents for each team, with the total number of agents being $N\hskip-1.4457pt=\hskip-1.4457ptN_{1}\hskip-1.4457pt+\hskip-1.4457ptN_{2}$. Let $X^{N_{1}}_{i,t}\in\mathcal{X}$ and $U^{N_{1}}_{i,t}\in\mathcal{U}$ represent the state and action of Blue agent $i\in[N_{1}]$ at time $t$. Here, $\mathcal{X}$ and $\mathcal{U}$ are the finite state and action spaces of the Blue team. Similarly, $Y^{N_{2}}_{j,t}\in\mathcal{Y}$ and $V^{N_{2}}_{j,t}\in\mathcal{V}$ denote the state and action of Red agent $j\in[N_{2}]$ at time $t$. The joint state-action variables for the Blue and Red teams are denoted as $(\mathbf{X}^{N_{1}}_{t},\mathbf{U}^{N_{1}}_{t})$ and $(\mathbf{Y}^{N_{2}}_{t},\mathbf{V}^{N_{2}}_{t})$, respectively. We denote the space of probability measures over a set $E$ as $\mathcal{P}({E})$. Below, $\mathrm{d}_{\mathrm{TV}}\big(\mu,\mu^{\prime}\big)$ represents the total variation between $\mu,\mu^{\prime}\in\mathcal{P}(E)$.
@@
We consider weakly-coupled dynamics where the dynamics of each individual agent is coupled with other agents through the EDs [Huang et al., 2006; Sanjari et al., 2022]. For Blue agent $i$, its stochastic transition is governed by the transition kernel $f_{t}:\mathcal{X}\times\mathcal{U}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\rightarrow\mathcal{P}(\mathcal{X})$:
@@
Lipschitz continuity is commonly assumed [Huang et al., 2006; Gu et al., 2021], and at minimum uniform continuity is required; see Cui et al. [2024] for counterexamples.
@@
We consider a mean-field sharing information structure [Arabneydi and Mahajan, 2015], where each agent’s decision depends on its own state and the two team EDs. We start with assuming full observation of mean-fields and later relax this assumption. Specifically, the Blue and Red agents seek to construct mixed Markov policies $\phi_{i,t}:\mathcal{U}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\to[0,1],~~\text{and}~~\psi_{j,t}:\mathcal{V}\times\mathcal{Y}\times\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y})\to[0,1]$, where the Blue policy $\phi_{i,t}(u|x_{i,t}^{N_{1}},\mu^{N_{1}}_{t},\nu^{N_{2}}_{t})$ dictates the probability that Blue agent $i$ selects action $u$ given its state $x_{i,t}^{N_{1}}$ and the observed/estimated team EDs $\mu^{N_{1}}_{t}$ and $\nu^{N_{2}}_{t}$. Note that each agent’s individual state is its private information.
@@ 2.2 Infinite-Population Solution
To reduce the complexity of team policy optimization domains in (4), the authors of Guan et al. [2024b] examined team behaviors under identical team policies at the infinite-population limit. It was shown that the team joint states can be represented using the team population distribution, which coincides with the state distribution of a typical agent referred to as the mean-fields ($\mu_{t}$ and $\nu_{t}$ for the Blue and Red teams, respectively). They also proved that MFs induced by identical team policies in an infinite-population game closely approximate the EDs induced by non-identical team policies in the corresponding finite-population game, which justifies the simplification of the optimization domain in (4) to identical team policies (also see Theorem 1). Furthermore, there is a one-to-one correspondence between infinite-population coordination policies $(\alpha,\beta)$ and local identical team policies $(\phi,\psi)\in\Phi\times\Psi$. The performance of $(\phi,\psi)$ in the equivalent zero-sum coordinator game is measured by
@@
where $\mu_{t}$ and $\nu_{t}$ follow a deterministic dynamics [Guan et al., 2024b] similar to the state distribution propagation of a controlled Markov chain. The worst-case performance of the Blue team in this infinite-population game is then given by the lower game value $\underline{J}_{\infty}^{*}(\mu_{0},\nu_{0})=\max_{\phi\in\Phi}\;\min_{\psi\in\Psi}~J_{\infty}^{\phi,\psi}(\mu_{0},\nu_{0})$, where the optimization domain is restricted to identical team policies. Guan et al. [2024b] establishes guarantees that identical team policies resulting from the solution of this equivalent zero-sum coordinator game are still $\epsilon$-optimal for the original max-min optimization problem in (4) where $\epsilon=\mathcal{O}(1/\sqrt{\underline{N}})$ and $\underline{N}=\min\{N_{1},N_{2}\}$.
The infinite-population limit of large-population games offers several theoretical advantages, such as representing the population by a typical agent and deterministic dynamics that reduce (3) to the non-stochastic optimization of (5). Previous works [Shao et al., 2024a; Perolat et al., 2021; Carmona et al., 2021] depend on infinite-population oracles to obtain mean-field trajectories $(\mu_{t},\nu_{t})$ in order to compute (5). This is rather unrealistic, since in practice, only finite-population simulations and local states $(\mathbf{x}^{N_{1}}_{t},\mathbf{y}^{N_{2}}_{t})$ with actions $(\mathbf{u}^{N_{1}}_{t},\mathbf{v}^{N_{2}}_{t})$ are available/observable. Moreover, a single coordinator policy $\alpha(\beta)$ defines a distribution over actions for each state conditioned on the mean-field, causing its dimensionality to scale with the joint state–action space (e.g., DDPG-MFTG), leading to high computational cost and degraded empirical performance (see Section 5). In summary, the infinite-population model is both impractical (due to oracle dependency) and computationally intractable (due to policy size). Thus, we turn to finite-population simulators and derive guarantees of optimality, scalability, and convergence of the policy gradient to the infinite-population ZS-MFTG.
The next result quantifies the level of suboptimality for the Blue team when it deploys the optimal identical policy learned directly from the solution of finite-population ZS-MFTG.
@@
Theorem 1 provides a principled justification for learning identical finite-population team policies in competitive–collaborative team games even when being exploited by non-identical opponent team strategies. Its motivation and proof build on the performance guarantees of the ZS-MFTG in the infinite-population limit, i.e., the coordinator game. Moreover, the error vanishes as $N_{1},N_{2}\to\infty$, thereby recovering the well-studied infinite-population MF formulation [Huang et al., 2006]. We detail this finite-population training paradigm in the next section.
@@ 3 Mean-Field Multi-Agent Proximal Policy Optimization
Motivated by Theorem 1, we present an algorithm to learn the finite-population optimal identical team policy. We build our algorithm based on the proximal policy optimization (PPO) framework due to its simplicity and effectiveness. While PPO has shown promising performance in cooperative tasks including MFC problems [Yu et al., 2022; Cui et al., 2024], its application in mixed competitive-collaborative scenarios is less studied. In the sequel, we introduce our key contribution: MF-MAPPO. We initialize two pairs of actor-critic networks, one for each team, deployed to learn the identical policy used by each team, see Figure 1(b). Specifically, we introduce a minimally-informed critic network by exploiting the MF information structure. The key point here is that we only require commonly accessible information for the critic network in order to learn the value function (Proposition 1). Further, the private information available to each agent only individually enters the actor during training. This results in neural networks that scale well with the number of agents. We present the team actor-critic networks from the Blue team’s perspective, and due to symmetry results extend naturally to the Red team.
@@
Shared-Team Actor. As discussed in earlier sections, the coordinator game is a useful theoretical construct but has limited practical value for real-world deployment since it relies on an infinite-population oracle for training and produces policies whose size scales poorly with the state-action space. We therefore directly train finite-population identical local policies, which preserve the mean-field structure while reducing complexity and improving tractability, with guarantees in Theorem 1. Not only is the approach computationally tractable in terms of the size of the policy, but is also more realistic in terms of sampling training data. We use a single actor network per team to learn identical team policies. The actor optimizes a PPO-based objective with a decaying entropy bonus [Schulman et al., 2017; Huang et al., 2022], which promotes exploration and stabilizes learning in mean-field settings [Cui and Koeppl, 2022; Guan et al., 2022]. Permutation invariance and identical team policies further allow a single replay buffer per team, reducing memory costs and simplifying experience collection. The PPO-based objective function of the Blue actor is given by:
@@
where, $g(\theta)=\frac{\phi_{\theta}(u|x,\mu,\nu)}{\phi_{\theta^{\text{old}}}(u|x,\mu,\nu)}$, $A_{k}$ is the generalized advantage function estimate [Schulman et al., 2018] and the tunable parameter $\omega$ weighs the contribution of the entropy term $S\big(\phi_{\theta}(x,\mu,\nu)\big)$ and decays as training progresses.